OpenStudy (anonymous):

2. The width w of a rectangular swimming pool is x + 4. The area A of the pool is 2x3 – 29x + 12. What is an expression for the length of the pool? (1 point)

OpenStudy (callisto):

length = area/ width = (2x^3 - 29x +12) / (x+4) = (x+4)(2x^2 -8x+3)/(x+4) =?